zunairah:

I need to use substitution and then graph the solution. I need help with the substitution part on this.

zunairah:

1 attachment
offbeatsoldier333334:

hmmmmmm

zunairah:

i dont get it

R3al:

Same

jhonyy9:

2x -4y = 4 2y = x-2 you need use the substitution method - do you know it how ? 2y = x-2 add to both sides 2 so will get expressed the x and this x value you need to substitute it in place of x inside the first equation

zunairah:

yea is know how to do substittuion

zunairah:

but the top equation confused me

zunairah:

since i need to plug it in for the y vale on the bottom equation

jhonyy9:

i m here to help you

jhonyy9:

no you express the x from second equation and you need substitute it in place of x inside the first equation

zunairah:

oh ok

jhonyy9:

on this way will get the y value what used you can calcule easy the x value

zunairah:

thank you!

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zunairah thank you! \(\color{#0cbb34}{\text{End of Quote}}\) np anytime - yw